OpenStudy (anonymous):

Using the given zero, find all other zeros of f(x). -2i is a zero of f(x) = x^4 - 32x^2 - 144

OpenStudy (ranga):

Complex roots always occur in pairs. If -2i is a root, then its complex conjugate, +2i will also be a root. That means, (x-2i) and (x+2i) are factors. (x-2i)(x+2i) = x^2 - (2i)^2 = (x^2 + 4) is a factor. Divide x^4 - 32x^2 - 144 by (x^2 + 4) It should divide evenly without any remainder. The quotient will be a quadratic which you can solve to find the other two roots.

OpenStudy (anonymous):

just to butt in i wouldn't actually divide using long division i would write \[x^4-32x-144=(x^2+4)(x^2+bx+c)\] and find \(b\) and \(c\)

OpenStudy (anonymous):

typo there, i meant \[x^4-32x^2-144=(x^2+4)(x^2+bx+c)\]

OpenStudy (anonymous):

Ok for ranga I got to x^4-176 x^2 - 576 and for satelite I got b=0 x=+- 2i c=-36

OpenStudy (anonymous):

I'm not sure how to go forward with either process

OpenStudy (anonymous):

Although 2i is one of the zeros in every solution I'm given!

OpenStudy (ranga):

Your long division is not correct. I will show you in the next reply. It is fairly straightforward here. You got b = 0 and c = -36 So the other factor is: x^2 +0x - 36 or x^2-36 which you can factor as: (x+6)(x-6) So the original function f(x) = (x-2i)(x+2i)(x-6)(x+6) The roots are: -2i, 2i, -6, 6.

OpenStudy (ranga):

x^2 - 36 ________________________ x^2 + 4 | x^4 - 32x^2 - 144 x^4 + 4x^2 -------------- -36x^2 - 144 -36x^2 - 144 ---------------- 0 0

OpenStudy (anonymous):

Well now I'm looking at what you did and I have no clue how I got my answer :/ Oh and duh, I had to factor 36! Thank you so much for your help :)